# On continuity of state-dependent utilities

### v1

@@ 3 Statement of the main results
This section is devoted to the proof of Theorem 3.1, which is quite involved and requires some preliminary groundwork.
@@
In the first part of the proof we lay the basis for the main argument by rephrasing some known results from WZ99 ; DM23 , obtaining in particular a functional that represents $\succeq$, which turns out to be additive on measurable sets. We provide some auxiliary lemmata: Lemma 4.6 proves that it is possible to construct an integrand function with the desired properties; this is based on Lemma 4.4 which firstly ensures the suitable measurability by a construction on rational numbers; Lemma 4.7 states that the integral representation is well defined and enjoys the desired properties; finally Lemma 4.8 shows continuity of $\mathbb{u}$. The concluding argument leads to Proposition 4.1 and illustrates how ultimately Theorem 3.1 follows. Consider a preference order $\succeq$ on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ satisfying (SM), (PC) and (ST). Suppose additionally that ${\mathcal{F}}$ contains at least three disjoint events $A_{1},A_{2},A_{3}$ such that $A_{i}\notin{\mathcal{N}}_{\succeq}$ for $i=1,2,3$. It is known (see WZ99 , Theorem 11) that $\succeq$ admits a numerical representation, i.e. there exists a functional $T:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ such that for any given $f,g\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$
@@
In particular WZ99 , Theorem 11, or CL06 , Theorem 4, imply that the Assumptions in DM23 , Theorem 5.4333Adopting ${\mathcal{F}}$ in place of ${\mathcal{G}}$, are met by the functional $T$ and we can therefore provide a further refinement of the representing functional. Specifically one can find a $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R},\;(A,f)\mapsto V_{A}(f)$ such that for $f_{1},f_{2}\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and $A\in{\mathcal{F}}$ we have $f_{2}\mathbf{1}_{A}\succeq f_{1}\mathbf{1}_{A}$ if and only if $V_{A}(f_{1})\leq V_{A}(f_{2})$. Furthermore, $V$ can be taken to satisfy the following properties (we refer the reader to the aforementioned literature for further details):
  1. (P1)
@@
  4. (P4)
    for every $A\in{\mathcal{F}}$ the functional $f\mapsto V_{A}(f)$, defined on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, is pointwise continuous, in that for any uniformly bounded [^4] Meaning $\sup_{n}\lVert f_{n}\rVert_{\infty}<+\infty$ sequence $(f_{n})_{n}\subset\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, $f_{n}(\omega)\to f(\omega)$ for any $\omega\in\Omega$, then $V_{A}(f_{n})\to V_{A}(f)$.
The following proposition provides a general representation result which stretches somewhat beyond the scope of the present work. A similar problem is in fact addressed within some relevant contributions in the stream of literature of AZ90 and (Buttazzo, , Ch. 2), under the name of Nonlinear Superposition Operator and exploiting the theory of Carathéodory functions. Another application of these representation results to subjective expected utility is presented in STANCA20 .
Proposition 4.1
[…]
@@
Remark 4.2
We emphasize that Proposition 4.1 provides a more general result than Theorem 3.1. Intuitively speaking, a functional that represents a preference relation as in WZ99 always satisfies the assumptions of the proposition, as will be explained more in detail at the end of this section.
Consider a functional $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ satisfying (P1), (P2), (P3) and (P4). For any $A\in{\mathcal{F}},f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ we denote
@@
and observe that, for all $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, $\mu_{f}$ is dominated by the probability $\mathbb{P}$ defined as in (P1). We now construct a special version of the state dependent utility, which enjoys the desired properties.
Definition 4.3
For any $q\in{\mathbb{Q}}$, we let $u(\cdot,q)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ be a version of the Radon-Nikodym derivative $\frac{\mathrm{d}\mu_{q}}{\mathrm{d}\mathbb{P}}$. In particular, we set $u(\omega,0)=0$ for all $\omega\in\Omega$ by virtue of (P1).
For any $A\in{\mathcal{F}}$ and $q\in\mathbb{Q}$, Definition 4.3 motivates the following notation
@@
which contradicts (P3) unless $\mathbb{P}(C)=0$. We conclude that ($\mathbb{P}(\{\omega\in\Omega:u(\omega,q_{1})<u(\omega,q_{2})\})=1$. By (countable) intersection over the rational numbers we obtain
@@
Since by (P3) and (P4) $V$ is strictly monotone and pointwise continuous we have that $\inf_{\tilde{q}>q}V_{A_{n}}(\tilde{q})=V_{A_{n}}(q)$, which in turn yields the desired contradiction. To conclude, since both (8) and (9) hold with probability $1$, we have
@@
Lemma 4.6
Let $u^{+}:\Omega\times\mathbb{R}\to\mathbb{R}$ be as in Definition 4.5. Then the following hold:
@@
    for every $\omega\in\Omega$ the function $u^{+}(\omega,\cdot)$ is RCLL555namely right continuous with left limits on $\mathbb{R}$;
Proof 2
@@
As to item (iii), fix any $x\in\mathbb{R}$ and consider $q_{1},q_{2}\in{\mathbb{Q}}$ such that $q_{1}\leq x\leq q_{2}$. Then by definition and item (i) $u(\omega,q_{1})\leq u^{+}(\omega,x)\leq u(\omega,q_{2})$ for any $\omega\in\Theta$. Since $\mathbb{P}(\Theta)=1$ and $u(\cdot,q_{1}),u(\cdot,q_{2})$ are versions of the Radon-Nikodym derivatives $\frac{d\mu_{q_{1}}}{\mathrm{d}\mathbb{P}},\frac{d\mu_{q_{2}}}{\mathrm{d}\mathbb{P}}\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$, integrability follows. Finally, observe that item (iv) holds trivially for $\omega\in\Omega\backslash\Theta$. Fix now $\omega\in\Theta$, we wish to show that for any $\varepsilon>0$ there exists $\delta$ such that $|u^{+}(\omega,x)-u^{+}(\omega,y)|<\varepsilon$ for any $y\in(x,x+\delta)$. Fix $\varepsilon>0$ and let $(q_{n})_{n}\subset{\mathbb{Q}}$ be such that $q_{n}\downarrow_{n}x$. Then there exists $\bar{n}\in\mathbb{N}$ such that, for all $n>\bar{n}$ we have $|u(\omega,q_{n})-\inf_{q>x}u(\omega,q)|=u(\omega,q_{n})-\inf_{q>x}u(\omega,q)<\varepsilon$. Recalling that, from item (i), $u(\omega,q_{n})=u^{+}(\omega,q_{n})$ and $\inf_{q>x}u(\omega,q)=u^{+}(\omega,x)$ by definition, monotonicity of $u^{+}$ implies that any $y\in(x,q_{\bar{n}})\subset\mathbb{R}$ satisfies $|u^{+}(\omega,x)-u^{+}(\omega,y)|<\varepsilon$. Existence of the left limit follows again from monotonicity proved in item (ii).
Lemma 4.7
Let $u^{+}:\Omega\times\mathbb{R}\to\mathbb{R}$ be as in Definition 4.5. Then the map
@@
Proof 3
Observe that, since $u^{+}(\omega,\cdot)$ is RCLL for every $\omega\in\Omega$, by Lemma A.1 we have that $T_{u^{+}}$ is well defined on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and monotone as $f(\omega)\leq g(\omega)$ implies $u^{+}(\omega,f(\omega))\leq u^{+}(\omega,g(\omega))$. We now show that $V_{\Omega}(f)=T_{u^{+}}(f)$ for every $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, then pointwise continuity of $T_{u^{+}}$ will follow from (P4). Fix any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, we can find two sequences of simple functions $\underline{f_{n}}=\sum_{A\in\underline{\pi}_{n}}\underline{x}^{n}_{A}\mathbf{1}_{A}$, $\overline{f_{n}}=\sum_{B\in\overline{\pi}_{n}}\overline{x}^{n}_{B}\mathbf{1}_{B}$, with $\overline{\pi}_{n},\underline{\pi}_{n}\subseteq{\mathcal{F}}$ finite partitions of $\Omega$, such that $\underline{f_{n}}\uparrow_{n}f,\overline{f_{n}}\downarrow_{n}f$. Moreover by standard arguments we can assume for any $n$ that $\underline{x}^{n}_{A},\overline{x}^{n}_{B}\in\mathbb{Q}$ for any $A\in\underline{\pi}_{n}$ and $B\in\overline{\pi}_{n}$. We have as a consequence of (7) and Lemma 4.6
@@
By (P4) we get the desired equality $V_{\Omega}(f)=T_{u^{+}}(f)$ for every $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$. Finally, we see that (P2) yields $V_{A}(f)=V_{\Omega}(f\mathbf{1}_{A})=T_{u^{+}}(f\mathbf{1}_{A})$, which concludes the proof.
Lemma 4.8
@@
Proof 4
We observe that, by Lemma 4.6, $u^{+}(\omega,\cdot)$ is strictly increasing, RCLL on $\mathbb{R}$ for all $\omega\in\Omega$ and $u^{+}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$. Moreover, as argued in Lemma 4.7, $T_{u^{+}}$ is pointwise continuous and in particular continuous from below. Consequently the assumptions of Corollary A.3 are verified and its application yields the function $\widehat{u^{+}}$ which we rename $\mathbb{u}:=\widehat{u^{+}}$. Finally, $\mathbb{u}$ satisfies all the requirements since
@@
Conclusions. We conclude the section by proving Proposition 4.1 and we show that Theorem 3.1 follows. Let $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ be a functional satisfying (P1), (P2), (P3) and (P4). Lemma 4.4 and 4.6 provide the construction of a suitable integrand function $u^{+}$ and Lemma 4.7 ensures that the map $f\mapsto\int_{\Omega}u^{+}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$ is well defined. Lemma 4.8 concludes the proof by showing the existence of the desired version $\mathbb{u}$ with the required properties. To prove Theorem 3.1, we exploit Proposition 4.1. Consider a preference $\succeq$ that is (SM), (PC) and (ST) and assume that ${\mathcal{F}}$ contains at least three disjoint events $A_{1},A_{2},A_{3}$ such that $A_{i}\notin{\mathcal{N}}_{\succeq}$ for $i=1,2,3$. As argued in the beginning of Section 4, the preference relation $\succeq$ induces a functional $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ that satisfies (P1), (P2), (P3) and (P4)  with $V_{\Omega}(\cdot)$ acting as numerical representation of the preference $\succeq$. Application of Proposition 4.1 provides a pair $(\mathbb{u},\mathbb{P})$ such that $V_{\Omega}(f)=\int_{\Omega}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$ for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$.
5 Proof of Theorem 3.2

### v2

@@ 5.2 Proof of Theorem 7
This section is devoted to the proof of Theorem 7, which is quite involved and requires some preliminary groundwork.
@@
In the first part of the proof we lay the basis for the main argument by rephrasing some known results from [24, 11], obtaining in particular a functional that represents $\succeq$, which turns out to be additive on measurable sets. We provide some auxiliary lemmata: Lemma 13 proves that it is possible to construct an integrand function with the desired properties; this is based on Lemma 12 which firstly ensures the suitable measurability by a construction on rational numbers; Lemma 14 states that the integral representation is well defined and enjoys the desired properties; finally Lemma 15 shows continuity of $\mathbb{u}$. The concluding argument leads to Proposition 11 and illustrates how ultimately Theorem 7 follows.
Consider a preference order $\succeq$ on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ satisfying (SM), (PC) and (ST). Suppose additionally that ${\mathcal{F}}$ contains at least three disjoint events $A_{1},A_{2},A_{3}$ such that $A_{i}\notin{\mathcal{N}}_{\succeq}$ for $i=1,2,3$. As proved in Proposition 9 (see also [24], Theorem 11), under these assumptions $\succeq$ admits a numerical representation, i.e. there exists a functional $T:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ such that for any given $f,g\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$
@@
In particular [24], Theorem 11, or [5], Theorem 4, imply that the Assumptions in [11], Theorem 5.4999Adopting ${\mathcal{F}}$ in place of ${\mathcal{G}}$, are met by the functional $T$ and we can therefore provide a further refinement of the representing functional. Specifically one can find a $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R},\;(A,f)\mapsto V_{A}(f)$ such that for $f_{1},f_{2}\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and $A\in{\mathcal{F}}$ we have $f_{2}\mathbf{1}_{A}\succeq f_{1}\mathbf{1}_{A}$ if and only if $V_{A}(f_{1})\leq V_{A}(f_{2})$. Furthermore, $V$ can be taken to satisfy the following properties (we refer the reader to the aforementioned literature for further details):
  1. (C1)
@@
  4. (C4)
    for every $A\in{\mathcal{F}}$ the functional $f\mapsto V_{A}(f)$, defined on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, is pointwise continuous, in that for any sequence $(f_{n})_{n}\subset\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, $\sup_{n}\lVert f_{n}\rVert_{\infty}<+\infty$, $f_{n}(\omega)\to f(\omega)$ for any $\omega\in\Omega$, then $V_{A}(f_{n})\to V_{A}(f)$.
The following proposition provides a general representation result which stretches somewhat beyond the scope of the present work. A similar problem is in fact addressed within some relevant contributions in the stream of literature of [2] and [4, Ch. 2], under the name of Nonlinear Superposition Operator and exploiting the theory of Carathéodory functions. Another application of these representation results to subjective expected utility is presented in [23].
Proposition 11.
Let $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ be a functional satisfying (C1), (C2), (C3)and (C4). Then there exists a ${\mathcal{F}}$-regular function $\mathbb{u}:\Omega\times\mathbb{R}\to\mathbb{R}$ with $\mathbb{u}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$, such that
@@
Consider a functional $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ satisfying (C1), (C2), (C3) and (C4). For any $A\in{\mathcal{F}},f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ we denote
@@
and observe that, for all $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, $\mu_{f}$ is dominated by the probability $\mathbb{P}$ defined as in (C1).
We now construct a special version of the state-dependent utility, which enjoys the desired properties.
Definition 5.
For any $q\in{\mathbb{Q}}$, we let $u(\cdot,q)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ be a version of the Radon-Nikodym derivative $\frac{\mathrm{d}\mu_{q}}{\mathrm{d}\mathbb{P}}$. In particular, we set $u(\omega,0)=0$ for all $\omega\in\Omega$ by virtue of (C1).
For any $A\in{\mathcal{F}}$ and $q\in\mathbb{Q}$, Definition 5 motivates the following notation
@@
which contradicts (C3) unless $\mathbb{P}(C)=0$. We conclude that ($\mathbb{P}(\{\omega\in\Omega:u(\omega,q_{1})<u(\omega,q_{2})\})=1$. By (countable) intersection over the rational numbers we obtain
@@
Since by (C3) and (C4) $V$ is strictly monotone and pointwise continuous we have that $\inf_{\tilde{q}>q}V_{A_{n}}(\tilde{q})=V_{A_{n}}(q)$, which in turn yields the desired contradiction. To conclude, since both (15) and (16) hold with probability $1$, we have
@@
Lemma 13.
Let $u^{+}:\Omega\times\mathbb{R}\to\mathbb{R}$ be as in Definition 6. Then the following hold:
@@
    for every $\omega\in\Omega$ the function $u^{+}(\omega,\cdot)$ is RCLL101010namely right continuous with left limits on $\mathbb{R}$;
Proof.
@@
As to item (iii), fix any $x\in\mathbb{R}$ and consider $q_{1},q_{2}\in{\mathbb{Q}}$ such that $q_{1}\leq x\leq q_{2}$. Then by definition and item (i) $u(\omega,q_{1})\leq u^{+}(\omega,x)\leq u(\omega,q_{2})$ for any $\omega\in\Theta$. Since $\mathbb{P}(\Theta)=1$ and $u(\cdot,q_{1}),u(\cdot,q_{2})$ are versions of the Radon-Nikodym derivatives $\frac{d\mu_{q_{1}}}{\mathrm{d}\mathbb{P}},\frac{d\mu_{q_{2}}}{\mathrm{d}\mathbb{P}}\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$, integrability follows.
Finally, observe that item (iv) holds trivially for $\omega\in\Omega\backslash\Theta$. Fix now $\omega\in\Theta$, we wish to show that for any $\varepsilon>0$ there exists $\delta$ such that $|u^{+}(\omega,x)-u^{+}(\omega,y)|<\varepsilon$ for any $y\in(x,x+\delta)$. Fix $\varepsilon>0$ and let $(q_{n})_{n}\subset{\mathbb{Q}}$ be such that $q_{n}\downarrow_{n}x$. Then there exists $\bar{n}\in\mathbb{N}$ such that, for all $n>\bar{n}$ we have $|u(\omega,q_{n})-\inf_{q>x}u(\omega,q)|=u(\omega,q_{n})-\inf_{q>x}u(\omega,q)<\varepsilon$. Recalling that, from item (i), $u(\omega,q_{n})=u^{+}(\omega,q_{n})$ and $\inf_{q>x}u(\omega,q)=u^{+}(\omega,x)$ by definition, monotonicity of $u^{+}$ implies that any $y\in(x,q_{\bar{n}})\subset\mathbb{R}$ satisfies $|u^{+}(\omega,x)-u^{+}(\omega,y)|<\varepsilon$. Existence of the left limit follows again from monotonicity proved in item (ii). ∎
Lemma 14.
Let $u^{+}:\Omega\times\mathbb{R}\to\mathbb{R}$ be as in Definition 6. Then the map
@@
Proof.
Observe that, since $u^{+}(\omega,\cdot)$ is RCLL for every $\omega\in\Omega$, by Lemma 19 we have that $T_{u^{+}}$ is well defined on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and monotone as $f(\omega)\leq g(\omega)$ implies $u^{+}(\omega,f(\omega))\leq u^{+}(\omega,g(\omega))$. We now show that $V_{\Omega}(f)=T_{u^{+}}(f)$ for every $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, then pointwise continuity of $T_{u^{+}}$ will follow from (C4). Fix any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, we can find two sequences of simple functions $\underline{f_{n}}=\sum_{A\in\underline{\pi}_{n}}\underline{x}^{n}_{A}\mathbf{1}_{A}$, $\overline{f_{n}}=\sum_{B\in\overline{\pi}_{n}}\overline{x}^{n}_{B}\mathbf{1}_{B}$, with $\overline{\pi}_{n},\underline{\pi}_{n}\subseteq{\mathcal{F}}$ finite partitions of $\Omega$, such that $\underline{f_{n}}\uparrow_{n}f,\overline{f_{n}}\downarrow_{n}f$. Moreover by standard arguments we can assume for any $n$ that $\underline{x}^{n}_{A},\overline{x}^{n}_{B}\in\mathbb{Q}$ for any $A\in\underline{\pi}_{n}$ and $B\in\overline{\pi}_{n}$. We have as a consequence of (14) and Lemma 13
@@
By (C4) we get the desired equality $V_{\Omega}(f)=T_{u^{+}}(f)$ for every $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$. Finally, we see that (C2) yields $V_{A}(f)=V_{\Omega}(f\mathbf{1}_{A})=T_{u^{+}}(f\mathbf{1}_{A})$, which concludes the proof. ∎
Lemma 15.
@@
Proof.
We observe that, by Lemma 13, $u^{+}(\omega,\cdot)$ is strictly increasing, RCLL on $\mathbb{R}$ for all $\omega\in\Omega$ and $u^{+}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$. Moreover, as argued in Lemma 14, $T_{u^{+}}$ is pointwise continuous and in particular continuous from below. Consequently the assumptions of Corollary 21 are verified and its application yields the function $\widehat{u^{+}}$ which we rename $\mathbb{u}:=\widehat{u^{+}}$. Finally, $\mathbb{u}$ satisfies all the requirements since
@@
Conclusions. We conclude the section by proving Proposition 11 and we show that Theorem 7 follows.
@@
Let $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ be a functional satisfying (C1), (C2), (C3) and (C4). Lemma 12 and 13 provide the construction of a suitable integrand function $u^{+}$ and Lemma 14 ensures that the map $f\mapsto\int_{\Omega}u^{+}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$ is well defined. Lemma 15 concludes the proof by showing the existence of the desired version $\mathbb{u}$ with the required properties.
To prove Theorem 7, we exploit Proposition 11. Consider a preference $\succeq$ that is (SM), (PC) and (ST) and assume that ${\mathcal{F}}$ contains at least three disjoint events $A_{1},A_{2},A_{3}$ such that $A_{i}\notin{\mathcal{N}}_{\succeq}$ for $i=1,2,3$. As argued in the beginning of Section 5.2, the preference relation $\succeq$ induces a functional $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ that satisfies (C1), (C2), (C3) and (C4)  with $V_{\Omega}(\cdot)$ acting as numerical representation of the preference $\succeq$. Application of Proposition 11 provides a pair $(\mathbb{u},\mathbb{P})$ such that $V_{\Omega}(f)=\int_{\Omega}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$ for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$.
5.3 Proof of Theorem 8